This is for solving praportions. What do you do if the cross multiply with the variable has a number more than one? cause I'm at this point in the problem x•15=3•1
Found 3 solutions by swincher4391, rfer, oberobic:
Answer by swincher4391(1107)

then you have 15x =3 And you solve for x. x = 3/15 = 1/5.

Answer by rfer(16322)

x/15=3/1
x=45

Answer by oberobic(2304)

If you're saying:
x*15 = 3*1
then there is no need to cross multiply.
Simply divide both sides by 15
x = 3/15
.
If you're saying:
x/15 = 3
then you can think of 3 as 3/1 and you can cross multiply
x = 45
.
Done.